# Unveiling Topological Structures from Language: A Comprehensive Survey of Topological Data Analysis Applications in NLP

### v3

@@ 7 Conclusion
Our world is currently experiencing an explosion of data and an equally explosion of computational techniques to process such data. Machine Learning (ML) is the most popular of these computational methods; however, while its benefits are numerous, it has a few limitations. The biggest of the limitations of ML is its inability to sufficiently process data that is high-dimensional, imbalanced, noisy, and scarce. Therefore, a small community of NLP researchers emerged to tackle this limitation by proposing using TDA to tackle difficult NLP tasks. These researchers employ two TDA techniques - Persistent Homology and Mapper to solve NLP tasks using theoretical and non-theoretical approaches. This yielded 95 papers which we comprehensively surveyed in this paper. Finally, we conclude that while the applications of TDA in NLP have greatly improved since 2013, there is still room for improvement, specifically in reducing the barrier to entry for non-TDA experts to apply it to their NLP tasks.
Acknowledgments
Adaku would like to thank Dr. Charlie Dagli for his encouragement, reading the paper drafts, and providing invaluable recommendations.
Appendix A Case Studies of TDA for NLP Tasks
We showcase the utility of TDA in non-trivial NLP tasks by applying persistent homology and mapper using the Theoretical and Non-theoretical approaches. See below:
A.1 Theoretical Approaches.
(Semantic) Topological Space.
For the analysis, we use a subset of the C4 dataset that has watermarked and non-watermarked LLM-generated texts [^4] https://huggingface.co/datasets/acmc/watermarked_c4_dataset. Next, we use the $ripser$ python package to extract the topological features with persistent homology and $textdescriptives$ Python package [^5] https://github.com/HLasse/TextDescriptives?tab=readme-ov-file to extract the linguistic features.
Topological “Shape” of Words.
Using the same C4 dataset, we create Mapper diagrams using the $Mapper$ Python package [^6] https://github.com/scikit-tda/kepler-mapper/tree/master.
A.2 Non-Theoretical Approaches
Model Interpretation and Analysis.
We evaluate the Euclidean and Wasserstein distances between the embeddings of English-French translation text pairs [^7] https://huggingface.co/datasets/aircrypto/English-French-Translations (using the first 500 text pairs). For the Euclidean distance, we compare the embeddings from the English model [^8] https://huggingface.co/sentence-transformers/stsb-xlm-r-multilingual and the multilingual model [^9] https://huggingface.co/sentence-transformers/msmarco-roberta-base-v2. To calculate the Wasserstein distance, we first extract topological features from the embeddings, represent them as persistent diagrams, and then calculate the distance. Euclidean and Wasserstein distance range is [0, $\infty$).
Appendix B List of papers
See list of Theoretical and Non-Theoretical applications of TDA in NLP in Tables 4 and 5, respectively.
Name Category Task TDA Technique
Chiang (2007) Topological Space (Semantic) Capturing semantic relationships within texts Persistent Homology
Karlgren et al. (2014) Topological Space (Semantic) Identify topical density of the space Mapper
Cavaliere et al. (2017) Topological Space (Semantic) Extracting extracts main concepts from text Persistent Homology
Wagner et al. (2012) Topological Space (Semantic) Analyzes document similarities Persistent Homology
Port et al. (2018) Topological Space (Syntactic) Analyzes syntactic parameters of different language families Persistent Homology
Port et al. (2022) Topological Space (Syntactic) Explaining linguistic structures with homoplasy phenomena Persistent Homology
Sami and Farrahi (2017) Topology of Topic Evolution Topic evolution within documents Persistent Homology
Draganov and Skiena (2024) Topological “Shape” of Words Investigate the “shape” of language phylogenies in the Indo-European language family Persistent Homology
Fitz (2022) Topological “Shape” of Words Capturing grammatical structure expressed by corpus using word manifold Persistent Homology
Fitz et al. (2024) Topological “Shape” of Words Measures the topological complexity of Transformer-based hidden representations Persistent Homology
Dong (2024) Topological “Shape” of Words Investigates the topological shapes of South American languages - the Nuclear-Macro-Jê (NMJ) and Quechuan families Persistent Homology
Bouazzaoui et al. (2021) Topological “Shape” of Words Investigates the similarity between the topological shapes of the Tifinagh and Phoenician scripts Persistent Homology
Table 4: Theoretical Applications of TDA in NLP.
Name Task Numerical Representation Learning Type Modality TDA Technique
SIFT Zhu (2013) child vs. adolescent writing detection TF-IDF Supervised Text Persistent Homology
Løvlie (2023) deepfake text detection TF-IDF, GloVe Supervised Text Persistent Homology
Huang (2022) election speech feature extraction TF-IDF Supervised Text Persistent Homology
Doshi and Zadrozny (2018); Shin (2019) movie genre classification TF-IDF Supervised Text Persistent Homology
Sovdat (2016) distinguishing between languages TF-IDF Supervised Text Persistent Homology
Savle et al. (2019) legal document entailment TF-IDF Supervised Text Persistent Homology
Michel et al. (2017) clustering and sentiment analysis TF-IDF, GloVe Supervised Text Persistent Homology
DoCollapse Guan et al. (2016) keyphrase extraction TF-IDF Unsupervised Text Persistent Homology
TOPOL Torres-Tramón et al. (2015) Twitter topic detection TF-IDF Supervised Text Persistent Homology
Kumar and Sarkar (2022) text summarization TF-IDF Unsupervised Text Persistent Homology
van Veen (2020) Propaganda tweet classification TF-IDF Unsupervised Text Mapper
Elyasi and Moghadam (2019) classify Persian poems TF-IDF Supervised Text Persistent Homology & Mapper
Effah (2017) age group categorization of lonely people TF-IDF Supervised Text Persistent Homology & Mapper
Maadarani and Hajra (2020) nursery rhyme classification from different continents - Australia, Asia, Africa, Europe, and North America TF-IDF Supervised Text Persistent Homology
Haghighatkhah et al. (2022) building document structure Pre-trained (Word2Vec) Unsupervised Text Persistent Homology
BERT+TDA Wu et al. (2022) contradiction detection Pre-trained (Word2Vec) Supervised Text Persistent Homology
Yessenbayev and Kozhirbayev (2022) speaker recognition & text processing Pre-trained (Word2Vec) Unsupervised Speech Persistent Homology
Yessenbayev and Kozhirbayev (2024) speaker recognition & text processing Pre-trained (Word2Vec) Unsupervised Speech Persistent Homology
Cornell (2020) building a topological search engine Pre-trained (Word2Vec) Unsupervised Text Mapper
Holmes (2020) document clustering and topic modeling tasks Pre-trained (Word2Vec) Supervised Text Mapper
Rawson et al. (2022) word sense induction and disambiguation Pre-trained (Word2Vec) Unsupervised Text Persistent Homology
Temčinas (2018) word sense induction and disambiguation Pre-trained (Word2Vec, GloVe) Unsupervised Text Persistent Homology
Feng et al. (2024) Geometry of textual data augmentation Pre-trained (Word2Vec) Supervised Text Persistent Homology
Tymochko et al. (2021) fraudulent paper detection Pre-trained (Word2Vec, GloVe, ElMo) Supervised Text Persistent Homology
Petri and Leitao (2020) disease epidemic prediction Pre-trained (Word2Vec) Supervised Text Persistent Homology
Tymochko et al. (2020) finding topological loops in logical statements Pre-trained (Word2Vec) Unsupervised Text Persistent Homology
Wright and Zheng (2020) distinguish subsets in data Pre-trained (Word2Vec) Unsupervised Text Persistent Homology
Paluzo Hidalgo et al. (2019) measuring the distance between the literary style of Spanish poets Pre-trained (Word2Vec) Supervised Text Persistent Homology
Gholizadeh et al. (2018) extract the topological signatures of novelists Pre-trained (GloVe) Supervised Text Persistent Homology
TIES Gholizadeh et al. (2020) document categorization & sentiment analysis Pre-trained (GloVe) Unsupervised Text Persistent Homology
Spannaus et al. (2024) phenotype prediction and news group categorization Pre-trained (GloVe) Unsupervised Text Mapper
Zadrozny (2021b) finding topological loops in logical statements Pre-trained (GloVe) Unsupervised Text Persistent Homology
Byers (2021) social anxiety detection Pre-trained (GloVe) Supervised Text Persistent Homology
Haim Meirom and Bobrowski (2022) compare cross-lingual sentence representations Pre-trained (GloVe) Unsupervised Text Persistent Homology
Zadrozny (2021a) investigates the manifestations of intelligence and understanding in neural networks Pre-trained (GloVe) Supervised Text Persistent Homology
Deng and Duzhin (2022) fake news detection Pre-trained (GloVe, BERT) Supervised Text Persistent Homology
Novak (2019) analyzing scholarly network Pre-trained (GloVe, BERT) Unsupervised Text Persistent Homology
Jakubowski et al. (2020) (1) polysemy word classification, and (2) word sense induction & disambiguation Pre-trained (FastText) Unsupervised Text Persistent Homology
Triki (2021); Shehu (2024) polysemy word classification Pre-trained (FastText) Supervised Text Persistent Homology
PHD Tulchinskii et al. (2024) deepfake text detection Pre-trained (Transformers - CLS) Supervised Text Persistent Homology
Short-PHD Wei et al. (2025) deepfake text detection (for short-text) Pre-trained (Transformers - CLS) Supervised Text Persistent Homology
Guilinger et al. (2025) deepfake text detection (for academic abstracts) Pre-trained (Transformers - CLS) Supervised Text Persistent Homology
Kushnareva et al. (2024) deepfake text detection Pre-trained (Transformers - CLS) Supervised Text Persistent Homology
Gourgoulias et al. (2024) class separability estimation Pre-trained (Transformers - CLS) Unsupervised Text Persistent Homology
TopoBERT Rathore et al. (2023) visually analyzing the fine-tuning process of a Transformer-based model Pre-trained (Transformers - CLS) Unsupervised Text Mapper
Lavery et al. (2024) fake news detection Pre-trained (Transformers - CLS) Supervised Text Persistent Homology
Das et al. (2021) classification of public speaking ratings from TED talks Pre-trained (Transformers - CLS) Supervised Text Persistent Homology
Byrne et al. (2022) topic modeling Pre-trained (Transformers - CLS) Unsupervised Text Mapper
TOPFORMER Uchendu et al. (2024) deepfake text detection Pre-trained (Transformers - Hidden) Supervised Text Persistent Homology
Garcia (2022) polysemy word classification Pre-trained (Transformers - Hidden) Supervised Text Mapper
Alexander and Wang (2023) Hate speech, Misinformation & Psychiatric disorder classification Pre-trained (Transformers - Hidden) Supervised Text Mapper
Persistent BERT Compression and Explainability (PBCE) Balderas et al. (2025) Model compression Pre-trained (Transformers - Hidden) Unsupervised Text Persistent Homology
Persistent Similarity Gardinazzi et al. (2024) Probing layers in LLMs Pre-trained (Transformers - Hidden) Unsupervised Text Persistent Homology
Sun and Nelson (2023) Correlation between sentence vectors Pre-trained (Word2Vec, Transformers - Hidden) Unsupervised Text Persistent Homology
Persistence Scoring Function Chauhan and Kaul (2022) captures the homology of the high-dimensional hidden representations Pre-trained (Transformers - Hidden) Unsupervised Text Persistent Homology
Topological Densification García-Castellanos et al. (2024) zero-shot model stitching Pre-trained (Transformers - Hidden) Unsupervised Text Persistent Homology
Ruppik et al. (2024) dialogue term extraction Pre-trained (Transformers - Hidden) Supervised Text Persistent Homology
Kudriashov et al. (2024) polypersonality Pre-trained (Transformers - Hidden) Unsupervised Text Persistent Homology
Bazarova et al. (2025) LLM Hallucination detection Pre-trained (Transformers - Attention) Supervised Text Persistent Homology
Kushnareva et al. (2021) deepfake text detection Pre-trained (Transformers - Attention) Supervised Text Persistent Homology
Cherniavskii et al. (2022); Proskurina et al. (2023); Jain et al. (2024) grammatical acceptability judgment Pre-trained (Transformers - Attention) Supervised Text Persistent Homology
Kostenok et al. (2023) Uncertainty estimation of model predictions Pre-trained (Transformers - Attention) Supervised Text Persistent Homology
Perez and Reinauer (2022) spam detection, grammatical acceptability judgment, and movie sentiment analysis Pre-trained (Transformers - Attention) Supervised Text Persistent Homology
Sakurai et al. (2025) Authorship attribution of Japanese texts Pre-trained (Transformers - Attention) Supervised Text Persistent Homology
Pollano et al. (2024) out-of-distribution detection Pre-trained (Transformers - Attention, CLS) Supervised Text Persistent Homology
Proskura and Zaytsev (2024) estimation of weights for ensembles of classification models Pre-trained (Transformers - Attention, CLS) Supervised Text Persistent Homology
Snopov and Golubinskiy (2024) vulnerability detection in code Pre-trained (Transformers - Attention) Supervised Text Persistent Homology
TopoHuBERT Tulchinskii et al. (2023) speaker recognition Pre-trained (Transformers - Attention) Supervised Speech Persistent Homology
Vukovic et al. (2022) dialogue term extraction Pre-trained (Transformers - Attention) Supervised Text Persistent Homology
Yurchuk and Gurnik (2023) Ukrainian tongue twisters classification Symbolic Representations Supervised Text Persistent Homology
Kovaliuk et al. (2024) Ukrainian tongue twisters classification Symbolic Representations Supervised Speech Persistent Homology
Bonafos et al. (2023) human vowel classification Multi-Modal Supervised Speech Persistent Homology
Bonafos et al. (2024) infant vocalization classification Multi-Modal Unsupervised Speech Persistent Homology
Gonzalez-Diaz et al. (2019); Paluzo-Hidalgo et al. (2022) emotion recognition Multi-Modal Supervised Speech Persistent Homology
Tlachac et al. (2020) depression detection Multi-Modal Supervised Speech Persistent Homology
Zhu et al. (2024) consonants recognition Multi-Modal Supervised Speech Persistent Homology
Sassone et al. (2022) music genre classification Multi-Modal Supervised Speech Persistent Homology
Vu et al. (2025) multi-modal adversarial robustness assessment Multi-Modal Supervised Image & Text Persistent Homology
Bergomi (2015) music classification Multi-Modal Supervised Speech Persistent Homology
Table 5: Non-theoretical applications of TDA in NLP

### v4

@@ 10 Conclusion
Our world is currently experiencing an explosion of data and an explosion of computational techniques to process such data. Machine Learning (ML) is the most popular of these computational methods; however, while its benefits are numerous, it has a few limitations. The biggest of the limitations of ML is its inability to sufficiently process data that is high-dimensional, imbalanced, noisy, and scarce. Therefore, a small community of NLP researchers emerged to tackle this limitation by proposing using TDA to tackle difficult NLP tasks. These researchers employ two TDA techniques - Persistent Homology and Mapper to solve NLP tasks using theoretical and non-theoretical approaches. This yielded 100 papers, which we comprehensively surveyed in this paper. Finally, we conclude that while the applications of TDA in NLP have improved greatly since 2013, there is still room for improvement, specifically in reducing the barrier to entry for non-TDA experts to apply it to their NLP tasks.
11 Ethical Statement
This survey highlights emerging applications of Topological Data Analysis (TDA) in Natural Language Processing. While our primary goal is to synthesize existing work, we recognize that several use cases carry important ethical considerations and dual-use risks. Topological methods can inadvertently expose latent sensitive attributes (e.g., dialect, health cues, authorship), enabling re-identification or profiling even when data is anonymized. Techniques discussed, such as watermarking, could also be repurposed to circumvent provenance systems. Applications in speech, emotion, and health-related domains further raise fairness, consent, and equity concerns, particularly for minority groups and low-resource languages. Therefore, we emphasize the need for bias and robustness audits, careful data governance and licensing, and privacy-preserving mechanisms when sharing derived features. Responsible release practices—such as restricting code that enables circumvention, conducting red-team evaluations, and requiring IRB or ethics review for clinical or surveillance-adjacent uses—are essential. Finally, given the computational demands of TDA pipelines, their environmental impact should be considered as well.